# PF-LHM: 3D Animatable Avatar Reconstruction from Pose-free Articulated Human Images

### v1

@@ 4.2 Qualitative Results
Table 5: Model Efficiency. We evaluate the training and inference efficiency of various backbones. The batch size is set to 1, the number of input views is equal to 16. ‘# Points’ refers to the number of geometric points, and ‘Time’ denotes the duration of a single iteration during both training and inference.
@@
LHM-0.5 B [31] 500 MB 80 K 26.33 s 54.97 GB 21.61 s 27.8 GB
LHM-0.7 B [31] 700 MB 80 K 50.23 s 62.27 GB 43.06 s 28.3 GB
LHM-1.0 B [31] 1000 MB 80 K 70.61 s 71.55 GB 64.29 s 29.7 GB
LHM-1.0 B [31] 1000 MB 160 K - $>$ 80 GB 159.19 s 35.8 GB
PF-LHM-S 500 MB 80 K 8.77 s 46.5 GB 7.35 s 17.0 GB
PF-LHM-M 700 MB 80 K 10.01 s 50.4 GB 9.03 s 18.3 GB
PF-LHM-L 1000 MB 80 K 11.13 s 55.1 GB 10.11 s 19.2 GB
PF-LHM-L 1000 MB 160 K 11.20 s 55.8 GB 10.64 s 20.0 GB
4.3 Ablation Study
Model Efficiency
Table 5 presents quantitative results that substantiate the efficiency of our model, tested on NVIDIA A100-80G hardware. The table clearly illustrates that our novel framework significantly outperforms the original LHM architecture, achieving notably reduced training times and lower memory consumption. Furthermore, during testing, our inference times are about 5$\sim$10 times faster than those of LHM.
Model Parameter Scalability
To verify the scalability of our PF-LHM, we train variant models with increasing parameter numbers by scaling the layer numbers. Table 6 compares performance across various model capacities. Our experiments indicate that increasing the number of model parameters correlates with improved performance. Figure 6 presents the comparison among PF-LHM-S, PF-LHM-M, and PF-LHM-L where the larger model achieves more accurate reconstruction. Also, Fig. 6 depicts the performance of different models based on varying input view counts, clearly indicating that our model is scalable and that performance improves with an increased number of input images.
Number of Query Points
Table 6 shows an ablation study analyzing the effect of varying the number of query points on public video datasets. As the number of query points increases from 40K to 80K, our model demonstrates improvements in PSNR, SSIM, and LPIPS by 0.562, 0.006, and 0.002, respectively. However, when the number of query points is increased further from 80K to 160K, we observe a slight gain in performance. Therefore, we set the number of query points to 80K to achieve an optimal balance between efficiency and model performance.
Table 6: Analysis of model parameters and 3D geometric point numbers.
[…]
PF-LHM-M 40 K 28.001 0.969 0.017 7.32 seconds 17 GB
PF-LHM-M 160 K 28.565 0.975 0.014 10.64 seconds 20 GB
PF-LHM-S 80 K 28.089 0.970 0.017 7.35 seconds 17 GB
PF-LHM-M 80 K 28.563 0.975 0.015 9.03 seconds 18 GB
PF-LHM-L 80 K 28.924 0.977 0.014 10.11 seconds 19 GB
5 Conclusion
We present PF-LHM, a novel feed-forward framework for rapid and high-fidelity 3D human avatar reconstruction from one or a few casually captured, pose-free images. Our approach introduces the Encoder-Decoder Point-Image Transformer (PIT), which enables efficient multimodal fusion between geometric point tokens and sparse multi-view image tokens through hierarchical attention. By leveraging the PIT architecture, our method achieves superior scalability, generalization, and reconstruction quality while significantly reducing runtime compared to optimization-based baselines. Extensive experiments across synthetic and real-world datasets demonstrate that PF-LHM effectively unifies single- and sparse-view reconstruction and supports realistic avatar animation using 3D Gaussian Splatting.
Appendix A Implementation Details
A.1 Details of Points Serialization
To trade the scalability and efficiency of our feed-forward framework, we leverage serialization to transform unstructured SMPL-X anchor points into structured data format. Following Point Transformer v3 (PTv3) [48], we mixture 4 patterns of serialization, including Z-order, Hibert, Trans Z-order and Trans Hibert, and apply random shuffle to the order of serialization patterns before each PTv3 block.
Figure 9: Detailed architecture of PTv3 Block [48].
A.2 Details of PTv3 Block
As illustrated in Fig. 9, our point attention architecture integrates the mechanism introduced in Point Transformer v3 [48]. This architecture employs patch-based self-attention to accelerate the forward pass. Furthermore, Grip Pooling is utilized to downsample the point cloud, enhancing the efficiency of point cloud self-attention. Importantly, to facilitate interaction among different patch groups, a shuffling order is implemented.
In our setup, the patch sizes are set to 4096, 2048, and 1024 for the 80K point clouds, decreasing in accordance with the downsampling process. For the 160K point clouds, the patch sizes are configured to 8192, 4096, and 2048.
Table 7: Ablation study on dataset scalability. The metrics are calculated based on 16-view image inputs.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
PF-LHM-L + Synthetic Data 26.148 0.952 0.027
PF-LHM-L + 10K Videos 27.456 0.970 0.018
PF-LHM-L + 100K Videos 28.312 0.975 0.017
PF-LHM-L + All 28.924 0.977 0.014
Only Synthetic 10K Videos 100K Videos 300K Videos GT
Figure 10: Ablation study on dataset scalability.
Ground Truth w/o Deformation w/ Deformation
Figure 11: The ablation study for posed-conditioned deformation. The red boxes highlight the visual difference mentioned above.
Input w/o $\mathcal{L}_{\text{dis}}$ w/ $\mathcal{L}_{\text{dis}}$ Human Mask
Figure 12: The ablation study for mask distribution loss.
Appendix B Experiments
B.1 Details of Testing Dataset
For our qualitative results, we utilize videos from Vid2Avatar [7], specifically those indexed as: 00000_random, 00020_Dance, 00069_Dance, exstrimalik, and Yuliang. Additionally, we draw on videos from Rec-MV [30], including the following indices: anran_dance_self_rotated, anran_purple, anran_skirt, self-rotate-leyang, and xiaolin. Moreover, we use the NeuMan [15] dataset with the indices: citron and seattle. And, sparse view images are randomly sampled from MVHumanNet [49], with the following indices: 101330, 101425, 101483, 103383, 103528, 104013, 203243, and 103594, along with videos captured using our smartphone.
B.2 More Ablation Study
Details of Dataset Scalability
To assess the scalability of our dataset, we perform controlled experiments utilizing stratified random subsets of 10K and 100K from the original training dataset of 300K videos. Table 7 demonstrates that relying solely on the synthetic dataset leads to poor model generalization. In contrast, incorporating an in-the-wild dataset significantly improves the model’s robustness and performance during real-world evaluations. Furthermore, increasing the size of the dataset siginificantly improves model results, although the gains in performance tend to diminish with larger dataset sizes. Figure 10 illustrates the findings of our ablation study regarding dataset scalability.
Mask Distribution Loss
Gaussian models tend to learn large-scale Gaussian primitives rather than large step offsets. To address this issue, we introduce a mask distribution regularization loss that encourages Gaussian primitives to learn offsets instead of concentrating on large-scale axes. As illustrated in Fig. 12, employing this loss enables the mean position of the Gaussian primitives to be distributed more evenly across the ground truth mask area, thereby preventing the model from learning large-scale axes of the Gaussian primitives.
Posed-Condition Deformation
Figure 11 illustrates the impact of using pose-conditioned deformation. The results clearly indicate that our deformation tokens, which are conditioned on SMPL-X parameters, effectively model non-rigid deformations of loose garments, in contrast to the case without pose conditioning.
B.3 More Results
@@
Figure 13 presents more animation results using eight input images. Our method enables high-fidelity reconstruction and animation of human avatars through the efficient Point-Image Transformer architecture, thereby demonstrating robust generalization capabilities and practical effectiveness.
Appendix C Limitations and Future Work
A primary limitation of our approach lies in its reliance on the SMPL-X template mesh for initializing geometric tokens. While this offers a strong structural prior, it can constrain reconstruction fidelity for subjects wearing loose or non-body-conforming garments, such as dresses, which deviate significantly from the SMPL-X topology. Additionally, due to the limited diversity of large motion poses in our training datasets, the model’s performance may degrade when encountering unseen or extreme poses. In future work, we plan to investigate more flexible garment representations and explore pose-independent anchor structures to better capture complex clothing dynamics and improve generalization to diverse motions.
@@ Appendix C Limitations and Future Work
Figure 13: More animation results of avatars created with 8-image inputs. Reference image is one of the input images.

### v2

@@ D.4 Training and Inference Efficiency
Table 12: End-to-end efficiency. Batch size=2, input views=8. # Points = geometric points; Time = per-iteration duration (training/inference).
@@
LHM-0.7B* [37] 700 M 40 K 81.22 s 80.3 GB 24.76 s 25.42 GB
700 M 160 K - OOM 110.3 s 36.71 GB
LHM++ Small 500 M 160 K 5.79 s 43.5 GB 1.58 s 6.98 GB
LHM++ Medium 700 M 160 K 6.35 s 48.7 GB 2.09 s 7.35 GB
LHM++ Large 900 M 160 K 7.25 s 52.6 GB 2.88 s 8.21 GB
D.5 More Quantitative Results
Table 3 of the main paper summarizes the averaged results on the public and smartphone-captured video sequences. Table 13 and Table 14 show the detailed results for each dataset.
Table 13: Comparison experiments with sparse-view input methods on public animation dataset.
views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++ M
2 21.718 0.929 0.125 3.6 m 21.049 0.937 0.086 3.8 m 26.371 0.959 0.045 8.5 m 26.900 0.964 0.022 1.08 s
4 21.782 0.926 0.126 6.0 m 22.614 0.943 0.065 5.6 m 26.906 0.960 0.044 15 m 27.031 0.965 0.022 1.40 s
8 22.308 0.928 0.120 10.6 m 23.238 0.947 0.061 9.0 m 27.319 0.962 0.040 32 m 27.346 0.965 0.021 2.09 s
16 23.026 0.936 0.101 18.8 m 23.427 0.949 0.061 15.7 m 27.456 0.962 0.040 1.2 h 27.521 0.966 0.020 3.66 s
64 23.561 0.944 0.077 54.0 m 23.574 0.949 0.055 1.1 h 27.896 0.962 0.037 4.4 h 27.557 0.966 0.020 17.94 s
Table 14: Comparison experiments with sparse-view input methods on smartphone-captured videos.
views InstantAvatar [17] GaussianAvatar [14] ExAvatar [29] LHM++ M
2 24.340 0.954 0.070 3.6 m 22.602 0.953 0.056 3.8 m 26.723 0.969 0.031 8.5 m 28.413 0.974 0.016 1.08 s
4 24.387 0.953 0.072 6.0 m 23.857 0.957 0.043 5.6 m 27.532 0.970 0.028 15 m 28.502 0.974 0.016 1.40 s
8 24.481 0.952 0.072 10.6 m 24.644 0.961 0.040 9.0 m 28.431 0.972 0.026 32 m 28.643 0.974 0.015 2.09 s
16 24.984 0.957 0.059 18.8 m 24.999 0.962 0.039 15.7 m 28.916 0.973 0.024 1.2 h 28.934 0.975 0.014 3.66 s
64 26.329 0.964 0.042 54.0 m 26.381 0.967 0.030 1.1 h 30.634 0.978 0.020 4.4 h 28.958 0.975 0.013 17.94 s
D.6 More Qualitative Results
Human Reconstruction in Canonical Space
Figure 15 demonstrates human reconstructions in canonical space from sparse views, showing that our model can reconstruct reasonable, high-quality avatars from sparse images input.
More Animation Rendering Results
@@ D.6 More Qualitative Results
Both Fig. 16 and Fig. 17 present more animation results using eight input images. Our method enables high-fidelity reconstruction and animation of human avatars through the efficient Point-Image Transformer architecture, thereby demonstrating robust generalization capabilities and practical effectiveness.
Appendix E Limitations and Future Work
A primary limitation of our approach is its dependence on the SMPL-X template mesh to initialize geometric tokens. While this provides a strong structural prior, it can limit reconstruction fidelity for subjects wearing loose or non-body-conforming garments, such as dresses and skirts. As shown in Figure 14, this manifests in failure cases where the model struggles to animate avatars wearing skirts during large lower-body movements. Furthermore, due to the limited diversity of large-motion poses in our training datasets, the model’s performance may decline when faced with unseen or extreme poses, which is also reflected in the failure cases presented. Additionally, since our model is primarily designed for sparse view input, we find that as the number of input views exceeds 16, the improvement in performance becomes marginal.
In future work, we plan to investigate more flexible garment representations and explore pose-independent anchor structures to better capture complex clothing dynamics, enhancing generalization to diverse motions. Moreover, our aim is to design a framework that is more suitable for dense view inputs.
Inputs Animation Results Inputs Animation Results
Figure 14: Failure Cases. our model struggles to animate avatars wearing skirts during large-amplitude lower‑body movements, and has difficulty handling inputs with extreme or challenging poses.
Reference Reconstruction Reference Reconstruction
Figure 15: Human reconstruction in canonical space from 8 input images. The reference image is one of the inputs.
@@ Appendix E Limitations and Future Work
Figure 16: More animation results of avatars created with 8-image inputs (Part I). Reference image is one of the input images.
Figure 17: More animation results of avatars created with 8-image inputs (Part II). Reference image is one of the input images.